OpenStudy (meilendurcer):

Use benchmarks to estimate each sum or difference. 17/18 - 3/7 A. about 1/4 B. about 1/2 C. about 0 D. about 1 i don't get it no matter how hard i try.

OpenStudy (across):

\[\frac ab\pm\frac cd=\frac{ad\pm bc}{bd}\]

OpenStudy (anonymous):

About 1

OpenStudy (across):

@RebelChi, you're off. ;P

OpenStudy (meilendurcer):

@across i have no idea how to do this

OpenStudy (anonymous):

Okay lol my teacher answered this one for me

OpenStudy (across):

@meilendurcer, the numbers \(17\) and \(18\) are close to each other (their difference is \(1\)), because of this, the ratio \(17/18\) is about \(1\). Moreover, the number \(7\) is over twice as large as the number \(3\). Because of this, the ratio \(3/7\) is a little less than half. If you subtract one-half from one, what do you get?

OpenStudy (meilendurcer):

1/4?

OpenStudy (across):

If I have a pineapple and I chew half of it off, how much will I have left?

OpenStudy (meilendurcer):

1/2

OpenStudy (across):

There you go.

OpenStudy (meilendurcer):

thank you